OpenStudy (anonymous):

Which lists the numbers in order from least to greatest? the mixed number one and two-fifths, the mixed number one and one-third, the decimal one and three tenths, the decimal one and twenty-five hundredths, the decimal one and two tenths the decimal one and two tenths, the mixed number one and two-fifths, the decimal one and twenty-five hundredths, the decimal one and three tenths, the mixed number one and one-third the decimal one and two tenths, the decimal one and twenty-five hundredths, the decimal one and three tenths, the mixed number one and one-third, the mixed number one an

OpenStudy (anonymous):

Convert every single one to a decimal and from there, you can organize the numbers.

OpenStudy (whpalmer4):

1 2/5 = 1 + 2/5 = 1 + 0.4 1 1/3 = 1 + 1/3 = 1 + 0.333333 1 3/10 = 1 + 3/10 = 1 + 0.3 1 25/100 = 1 + 25/100 = 1 + 0.25 1 2/10 = 1 + 2/10 = 1 + 0.2 put them in order and figure out which answer choice orders them from least to greatest